# TimeRecipe: A Time-Series Forecasting Recipe via Benchmarking Module Level Effectiveness

### v1

@@ 4.1.1 TimeRecipe Yields Model Architectures Outperforming SOTA
More importantly, this phenomenon is not a rare case: it holds in 92 out of 102 evaluated scenarios, detailed in Appendix C.2, suggesting that in over 90% of cases, TimeRecipe can identify a model architecture that surpasses the SOTA. On average, the best existing approaches lag behind the top-performing configurations identified by TimeRecipe by 13.66 ranking positions. By exhaustively exploring the design space, TimeRecipe achieves an average forecasting error reduction of 5.4% compared to the best existing methods.
@@ 4.1.2 TimeRecipe Uncovers Modules Effectiveness Linking to Data Properties
To address this gap and answer our central research question: Which modules and model designs are most effective under specific time-series forecasting scenarios? As a first step, we establish a taxonomy of time-series data characteristics. This taxonomy includes key properties such as seasonality, trend, stationarity, transition, shifting, and correlation qiu2024tfb , as detailed in Appendix B.3. In addition to these intrinsic data properties, we further categorize forecasting tasks based on input structure—namely, whether the data is univariate or multivariate—as well as the number of input features (N-Feature) and the horizon-to-lookback ratio (HL-Ratio), which may also influence the effectiveness of specific module configurations.
@@
Building on these data characteristics, we conduct a correlation analysis to explore the relationship between dataset properties and the effectiveness of various module designs. Specifically, we employ statistical hypothesis testing via t-tests to assess whether particular module configurations yield significantly improved performance under specific data conditions. This analysis directly supports our primary research goal of identifying the most appropriate design choices for diverse forecasting scenarios. A complete list of all statistically significant correlations (i.e., $p$-value $\leq 0.05$) is provided in Table 3, Appendix C.1. Through the analysis, we observe patterns among the statistically significant module-level effectiveness as follows:
Pre-processing. Time-series pre-processing techniques, such as instance normalization and series decomposition, consistently enhance ranking performance under complex conditions. These include high correlation, transition, seasonality, and shifting, as well as high HL-Ratio (i.e., long-term forecasting) and low stationarity. This suggests that pre-processing stabilizes input representations, especially for irregular time series.
Fusion. Temporal fusion is broadly effective in different settings. It improves performance when the n-feature is low, and under high HL-Ratio, correlation, transition, shifting, seasonality, and trend, highlighting the importance of capturing contextual temporal patterns in diverse forecasting scenarios.
Embedding Strategies. Embedding choices are more dependent on data properties. Invert embedding is favored when the n-feature, correlation, or transition is high. Token embedding is more effective under low HL-Ratio, trend, or stationarity. Patch embedding performs well in low-trend regimes. In relatively simple settings (e.g., low-trend univariate series), no embedding can also be beneficial.
Feedforward Modules. Transformer-based feedforward layers are favorable when correlation, seasonality, or transition is high, reflecting their strength in capturing global dependencies. Conversely, RNN-based feedforward layers perform better when n-feature count, HL-Ratio, shifting, trend, or stationarity is low, suggesting their suitability for simpler or more sequential patterns.
These findings suggest that architectural decisions should be tailored to the underlying data characteristics, with certain configurations consistently outperforming others under specific conditions.
@@ 4.2 Practical Implications from the Insights
Given the insights discussed above, a natural question arises: how can these findings be leveraged to inform better model architecture design for a given time-series dataset? To address this, TimeRecipe integrates a training-free model selection mechanism based on a LightGBM ke2017lightgbm regression model. Specifically, we train a regression model that maps both dataset characteristics and model configurations to their associated rank scores benchmarked by TimeRecipe. For a new forecasting task, we first compute the relevant data characteristics and then estimate the rank scores for a range of candidate configurations. The configuration with the lowest predicted rank score will be selected.
To assess the effectiveness of this training-free selection strategy, we evaluate it on two forecasting scenarios: (i) an in-distribution case involving short-term multivariate forecasting on the ETTh1 dataset (which is used in the TimeRecipe, though this specific setting is not benchmarked), and (ii) an out-of-distribution case using a new univariate unemployment (social) forecasting dataset introduced in Time-MMD liu2024timemmd , which is not part of the original TimeRecipe benchmark.
The results, presented in Table 2, demonstrate that even with a simple tree-based approach, TimeRecipe is capable of selecting models that are closer to the globally optimal architecture than the existing best-performing model. These findings underscore the potential of deploying TimeRecipe in real-world model selection scenarios, where it is able to make model selections that can surpass the SOTA performance without the need for training.
ETTh1_24_M
IN SD Fusion Embed FF-Type Rank MSE MAE
TimeRecipe Best Feature Patch MLP 1.5 0.2963 - 0.3467 -
Top-3 Selection Feature Patch MLP 1.5 0.2963 0.0% 0.3467 0.0%
Temporal Patch MLP 20.5 0.3096 -4.3% 0.3586 -3.4%
Temporal Patch Trans. 4.0 0.2988 -0.9% 0.3520 -1.5%
Selection better than existing best: Yes (Selection 1&3)
@@
IN SD Fusion Embed FF-Type Rank MSE MAE
@@
Table 2: Comparison of model configurations selected by our method against the best-performing existing model and the global best results found through exhaustive design space search. On both ETTh1_24_M and Social_12_S, at least one of our top-3 selections consistently outperforms the existing best and closely approaches the optimal configuration. This demonstrates the effectiveness of our training-free model selection method, even when using a simple tree-based approach.
@@ 4.3 Implications for Future Research
Implications for Designing Advanced Time-Series Forecasting Models. TimeRecipe indicates that no single existing architecture consistently outperforms others across all time-series forecasting scenarios, even with taking recent SOTA into consideration shi2024time ; wang2024timemixer++ ; zhong2025time ; kong2025unlocking . However, rather than interpreting this as a limitation that precludes the development of broadly effective models, we argue that this insight points to the potential of hybrid architectures. Future SOTA models may benefit from integrating multiple architectural components tailored to different data properties. For example, ni2025timedistill has shown that some temporal patterns are better captured by Transformer-based models, while others are more effectively modeled by MLPs. Extending this intuition, TimeRecipe aims to reveal that similar complementarities exist across a wide range of design dimensions, such as normalization strategies, embedding schemes (e.g., patch vs. token), or architectural combinations. These findings suggest that next-generation forecasting systems, including foundation models, which are often derived from supervised learning approaches, could possibly achieve broader robustness and improved generalization by dynamically leveraging hybrid designs tailored to specific forecasting scenarios and pattern types.
Advancing AutoML for Time-Series Forecasting. Another important contribution of TimeRecipe lies in its potential to advance AutoML research in time-series forecasting. While earlier literature has explored AutoML for time series, these studies have predominantly focused on model-level searches combined with hyperparameter and learning scheme optimization alsharef2022time ; alsharef2022review ; shchur2023autogluon ; westergaard2024time . With the canonical architecture introduced by TimeRecipe, future AutoML for time-series forecasting may enable finer-grained, module-level search across different components of the forecasting pipeline. This approach allows AutoML systems to explore a significantly broader and more flexible design space, in conjunction with traditional hyperparameter and learning scheme tuning. As a result, TimeRecipe potentially benefits more granular and effective AutoML in time-series applications.
@@ Appendix A Related Works
Supervised Time-Series Forecasting. Time-series forecasting, predicting future values based on historical observations, is a long-standing challenge across numerous domains. Early approaches have been dominated by statistical methods. Autoregressive Integrated Moving Average (ARIMA) models [8] and Exponential Smoothing (ETS) techniques [25] have become foundational, modeling temporal dependencies and trends through established statistical principles. These methods often assume linearity or specific data structures and can struggle with complex, non-linear patterns present in many real-world datasets.
@@
As a transformative architecture from deep learning, the introduction of the Transformer architecture [65], with its powerful self-attention mechanism, has marked a pivotal moment for time-series forecasting. Early efforts have focused on enhancing the efficiency and effectiveness of Transformers for long sequential data, a common characteristic of time series. Models like Reformer [29] have introduced techniques to reduce computational complexity, while Informer [86], Autoformer [75], Pyraformer [40], and FEDformer [87] have explored specialized attention mechanisms (e.g., sparse attention, auto-correlation, pyramidal attention) and decomposition strategies (e.g., trend-seasonal decomposition, frequency domain processing) to better capture temporal dependencies in long sequences. Further refinements have addressed specific challenges inherent in time-series data [35]. For instance, Non-stationary Transformers [44] have incorporated mechanisms to handle distribution shifts over time, ETSformer [72] has integrated principles from classical exponential smoothing, and Crossformer [82] has focused on modeling dependencies across different variates in multivariate settings. PatchTST [54] has proposed segmenting time series into patches, treating them as tokens, which proves highly effective for long-term forecasting. More recently, iTransformer [42] has inverted the roles of embedding and attention layers, achieving strong results.
@@
Despite the success of Transformer variants, recent research has spurred a debate regarding their necessity, demonstrating that simpler architectures, particularly those based on Multi-Layer Perceptrons (MLPs) or even linear layers, can achieve competitive or superior performance on many benchmarks. DLinear [79] has proposed a simple linear model with decomposition, challenging the complexity of contemporary Transformers. Subsequently, various MLP-based models have emerged, often emphasizing efficiency and specialized designs. LightTS [81] has utilized sampling-oriented MLP structures, TSMixer [13] has employed an all-MLP architecture with mixing across time steps and features. TimeMixer [69] and TimeMixer++ [68] have further reflected on the competent duties of Transformer components and have repurposed the Transformer architecture.
@@
Beyond Transformers and MLPs, other architectural paradigms continue to be explored. Convolutional approaches, such as MICN [66] using multi-scale local and global context modeling, SCINet [39] employing sample convolution and interaction and TimesNet [74] modeling temporal 2D variations, offer alternative ways to capture temporal features. Recurrent architectures are also being revisited, exhibiting strong performance for forecasting [30]. Recent explorations also include leveraging pretrained language models for forecasting tasks [26, 63, 36]. This diverse landscape, i.e., various model architectures, highlights the necessity and urgency of this time-series forecasting recipe at the model level, which is developed in this paper.
@@
Time-Series Foundation Models. Time-series foundation models, pre-trained on vast amounts of diverse time-series data, aim to zero-shot adapt to unseen time-series datasets. Existing methods largely follow the architectural design of foundation language models. For instance, Chronos [5] is based on the T5 [57] architecture, TimesFM [16] employs the decoder-only Transformer architecture, Lag-Llama [59] explicitly leverages the Llama architecture. Recent advancements [41, 62] address computational scaling while enhancing capacity by introducing mixture of experts (MoE) [11] techniques from foundation language models. We envision that the model-level time-series forecasting recipe proposed in this paper provides key insights for the architectural design of the next generation of foundation time-series models.
@@
Time-Series Benchmarks. Robust and standardized benchmarks are essential for evaluating the performance of time-series forecasting models, enabling fair comparison and reproducible research. Historically, the M4 and M5 competitions [49, 50] have served as crucial benchmarks, providing large collections of diverse time series primarily in business, demographic, and economic domains, along with rigorous evaluation protocols [49, 50]. Recent efforts focus on providing unified benchmarking frameworks and toolkits. TFB [56], for instance, introduces a scalable suite covering multiple domains and methods, emphasizing reproducibility and systematic, aspect-based analysis. Furthermore, BasicTS [34] focuses on benchmarking multivariate time series forecasting, ProbTS [80] benchmarks both point and distributional forecasting, and GIFT-Eval [2] is designed for benchmarking foundation time-series models. Recently, ReC4TS [38] benchmarks how different reasoning strategies enhance zero-shot time-series forecasting. In addition, Time-MMD [37] and CiK [6] are designed for benchmarking multimodal time-series forecasting. Different from all existing time-series benchmarks, our work uniquely focuses on evaluating the effectiveness of individual module-level design choices for time-series forecasting. To the best of our knowledge, this is the first benchmark to systematically address this underexplored yet critical aspect of model design.
@@
Auto-ML in Time-Series. Automated Machine Learning (AutoML) aims to automate the end-to-end process of applying machine learning, enhancing accessibility and efficiency. AutoML research in time series can be categorized as: (1) automated feature engineering, which generates relevant temporal features like lags and rolling statistics [12, 15]; (2) automated model selection, searching across diverse model families from statistical methods (e.g., ARIMA) to machine learning (e.g., boosted trees) and deep learning architectures (e.g., LSTMs, Transformers) [78, 1, 61]; (3) hyperparameter optimization, using techniques like Bayesian optimization or evolutionary algorithms to tune model configurations [73, 22]; (4) neural architecture search, which specifically automates the design of deep learning model structures suitable for capturing complex temporal patterns [58, 76].

### v2

@@ 4.1.1 TimeRecipe Yields Model Architectures Outperforming SOTA
[…]
@@ 4.1.2 TimeRecipe Uncovers Modules Effectiveness Linking to Data Properties
To address this gap and answer our central research question: Which modules and model designs are most effective under specific time-series forecasting scenarios? As a first step, we establish a taxonomy of time-series data characteristics. This taxonomy includes key properties such as seasonality, trend, stationarity, transition, shifting, and correlation (Qiu et al., 2024), as detailed in Appendix B.3. In addition to these intrinsic data properties, we further categorize forecasting tasks based on input structure, namely, whether the data is univariate or multivariate, as well as the number of input features (N-Feature) and the horizon-to-lookback ratio (HL-Ratio), which may also influence the effectiveness of specific module configurations.
@@
Building on these data characteristics, we conduct a correlation analysis to explore the relationship between dataset properties and the effectiveness of various module designs. Through the analysis, we observe patterns among the statistically significant module-level effectiveness, as shown in Table 2. Specifically, we employ statistical hypothesis testing via t-tests to assess whether particular module configurations yield significantly improved performance under specific data conditions. This analysis directly supports our primary research goal of identifying the most appropriate design choices for diverse forecasting scenarios. A complete list of all statistically significant correlations (i.e., $p$-value $\leq 0.05$) is provided in Table 4, Appendix C.1.
Setting Choice Condition
Multivariate Instance Norm.
when n-feature is low hl-ratio is high correlation is high
shifting is high trend is high seasonality is low
Series Decomp.
when shifting is low
Temporal Fusion
when n-feature is low hl-ratio is high correlation is low
transition is high shifting is high trend is high
seasonality is low
Invert Embed.
when n-feature is high correlation is high transition is low
shifting is low
Token Embed.
when hl-ratio is low trend is low seasonality is high
stationarity is low
Patch Embed.
when trend is high
RNN Arch.
when n-feature is high hl-ratio is low shifting is low
trend is low stationarity is low
Transformer Arch.
when correlation is high transition is low seasonality is high
Univariate Series Decomp.
when transition is high shifting is high trend is high
stationarity is high
Temporal Fusion
when hl-ratio is high shifting is high trend is high
seasonality is low
Non-Embed.
when trend is high
Transformer Arch.
when seasonality is high
Table 2: Module choices under specific data property conditions.
These findings provide a more granular view of module-level effectiveness, some of which align with well-established knowledge. For example, instance normalization, which is specifically designed to handle distribution shifts, is most beneficial when shifts are large or seasonality is low. Similarly, RNN architectures are more flexible when the HL-ratio is low, as errors tend to accumulate over longer horizons. Furthermore, our results highlight directions for future in-depth studies. For instance, while this work provides empirical insights, further investigation is warranted to understand how time series decomposition interacts with shifting patterns: being more effective under less shifting in multivariate setups and under more shifting in univariate setups. Overall, these results suggest that architectural choices should be carefully aligned with the characteristics of the underlying data, with certain configurations consistently outperforming others under specific conditions.
@@ 4.2 Practical Implications from the Insights
Given the insights discussed above, a natural question arises: how can these findings be leveraged to inform better model architecture design for a given time-series dataset? To address this, TimeRecipe integrates a training-free model selection mechanism based on a LightGBM (Ke et al., 2017) regression model. Specifically, we train a regression model that maps both dataset characteristics and model configurations to their associated rank scores benchmarked by TimeRecipe. For a new forecasting task, we first compute the relevant data characteristics and then estimate the rank scores for a range of candidate configurations. The configuration with the lowest predicted rank score will be selected.
We evaluate the model selection toolkit on two forecasting scenarios: (i) an in-distribution case involving short-term multivariate forecasting on the ETTh1 dataset (which is used in the TimeRecipe, though this specific setting is not benchmarked), and (ii) three out-of-distribution cases using a new univariate unemployment forecasting dataset introduced in Time-MMD (Liu et al., 2024a), which is not part of the original TimeRecipe benchmark. The results, presented in Table 3, and Table 5 in Appendix C.2, demonstrate that even with a simple tree-based approach, TimeRecipe is capable of selecting models that are closer to the globally optimal architecture than the existing best-performing model. These findings underscore the potential of deploying TimeRecipe in real-world model selection scenarios, where it is able to make model selections that can surpass the SOTA performance without the need for training.
@@
IN SD Fusion Embed $F$-Type Rank MSE MAE
@@
Table 3: Comparison of model configurations selected by our method against the best-performing existing model and the global best results found through exhaustive design space search. At least one of our top-3 selections consistently outperforms the existing best and closely approaches the optimal configuration on Social_12_S (see Table 5, Appendix C.2, for more and similar observations). This demonstrates the value of TimeRecipe, even when using a simple tree-based approach.
@@ 4.3 Implications for Future Research
Implications for Designing Advanced Time-Series Forecasting Models. TimeRecipe indicates that no single existing architecture consistently outperforms others across all time-series forecasting scenarios, even with taking recent SOTA into consideration (Shi et al., 2024; Wang et al., 2025; Zhong et al., 2025; Kong et al., 2025). However, rather than interpreting this as a limitation that precludes the development of broadly effective models, we argue that this insight points to the potential of hybrid architectures. Future SOTA models may benefit from integrating multiple architectural components tailored to different data properties. For example, (Ni et al., 2025) has shown that some temporal patterns are better captured by Transformer-based models, while others are more effectively modeled by MLPs. Extending this intuition, TimeRecipe aims to reveal that similar complementarities exist across a wide range of design dimensions, such as normalization strategies, embedding schemes (e.g., patch vs. token), or architectural combinations. These findings suggest that next-generation forecasting systems, including foundation models, which are often derived from supervised learning approaches, could possibly achieve broader robustness and improved generalization by dynamically leveraging hybrid designs tailored to specific forecasting scenarios and pattern types.
Advancing AutoML for Time-Series Forecasting. Another important contribution of TimeRecipe lies in its potential to advance AutoML studies in time-series forecasting. Earlier studies have mainly focused on model-level searches combined with hyperparameter and learning scheme optimization (Alsharef et al., 2022b; a; Shchur et al., 2023; Westergaard et al., 2024). With the canonical architecture introduced by TimeRecipe, future AutoML for time-series forecasting may enable finer-grained, module-level search across different components of the forecasting pipeline. This approach allows AutoML systems to explore a significantly broader and more flexible design space, in conjunction with traditional hyperparameter and learning scheme tuning. As a result, TimeRecipe potentially benefits more granular and effective AutoML in time-series applications.
@@ Appendix A Related Works
Supervised Time-Series Forecasting. Time-series forecasting, predicting future values based on historical observations, is a long-standing challenge across numerous domains. Early approaches have been dominated by statistical methods. Autoregressive Integrated Moving Average (ARIMA) models (Box, 2013) and Exponential Smoothing (ETS) techniques (Hyndman et al., 2008) have become foundational, modeling temporal dependencies and trends through established statistical principles. These methods often assume linearity or specific data structures and can struggle with complex, non-linear patterns present in many real-world datasets.
@@
As a transformative architecture from deep learning, the introduction of the Transformer architecture (Vaswani et al., 2017), with its powerful self-attention mechanism, has marked a pivotal moment for time-series forecasting. Early efforts have focused on enhancing the efficiency and effectiveness of Transformers for long sequential data, a common characteristic of time series. Models like Reformer (Kitaev et al., 2020) have introduced techniques to reduce computational complexity, while Informer (Zhou et al., 2021), Autoformer (Wu et al., 2021), Pyraformer (Liu et al., 2022b), and FEDformer (Zhou et al., 2022) have explored specialized attention mechanisms (e.g., sparse attention, auto-correlation, pyramidal attention) and decomposition strategies (e.g., trend-seasonal decomposition, frequency domain processing) to better capture temporal dependencies in long sequences. Further refinements have addressed specific challenges inherent in time-series data (Liu et al., 2024a). For instance, Non-stationary Transformers (Liu et al., 2022c) have incorporated mechanisms to handle distribution shifts over time, ETSformer (Woo et al., 2022) has integrated principles from classical exponential smoothing, and Crossformer (Zhang & Yan, 2023) has focused on modeling dependencies across different variates in multivariate settings. PatchTST (Nie et al., ) has proposed segmenting time series into patches, treating them as tokens, which proves highly effective for long-term forecasting. More recently, iTransformer (Liu et al., ) has inverted the roles of embedding and attention layers, achieving strong results.
@@
Despite the success of Transformer variants, recent research has spurred a debate regarding their necessity, demonstrating that simpler architectures, particularly those based on Multi-Layer Perceptrons (MLPs) or even linear layers, can achieve competitive or superior performance on many benchmarks. DLinear (Zeng et al., 2023) has proposed a simple linear model with decomposition, challenging the complexity of contemporary Transformers. Subsequently, various MLP-based models have emerged, often emphasizing efficiency and specialized designs. LightTS (Zhang & Yan, 2022) has utilized sampling-oriented MLP structures, TSMixer (Chen et al., 2023) has employed an all-MLP architecture with mixing across time steps and features. TimeMixer (Wang et al., 2024a) and TimeMixer++ (Wang et al., 2025) have further reflected on the competent duties of Transformer components and have repurposed the Transformer architecture.
@@
Beyond Transformers and MLPs, other architectural paradigms continue to be explored. Convolutional approaches, such as MICN (Wang et al., 2023) using multi-scale local and global context modeling, SCINet (Liu et al., 2022a) employing sample convolution and interaction and TimesNet (Wu et al., ) modeling temporal 2D variations, offer alternative ways to capture temporal features. Recurrent architectures are also being revisited, exhibiting strong performance for forecasting (Kong et al., 2025). Recent explorations also include leveraging pretrained language models for forecasting tasks (Jin et al., ; Tan et al., 2024; Liu et al., 2025a). This diverse landscape, i.e., various model architectures, highlights the necessity and urgency of this time-series forecasting recipe at the model level, which is developed in this paper.
@@
Time-Series Foundation Models. Time-series foundation models, pre-trained on vast amounts of diverse time-series data, aim to zero-shot adapt to unseen time-series datasets. Existing methods largely follow the architectural design of foundation language models. For instance, Chronos (Ansari et al., ) is based on the T5 (Raffel et al., 2020) architecture, TimesFM (Das et al., ) employs the decoder-only Transformer architecture, Lag-Llama (Rasul et al., ) explicitly leverages the Llama architecture. Recent advancements (Liu et al., 2024b; Shi et al., 2024) address computational scaling while enhancing capacity by introducing mixture of experts (MoE) (Cai et al., 2025) techniques from foundation language models. We envision that the model-level time-series forecasting recipe proposed in this paper provides key insights for the architectural design of the next generation of foundation time-series models.
@@
Time-Series Benchmarks. Robust and standardized benchmarks are essential for evaluating the performance of time-series forecasting models, enabling fair comparison and reproducible research. Historically, the M4 and M5 competitions (Makridakis et al., 2020; 2022) have served as crucial benchmarks, providing large collections of diverse time series primarily in business, demographic, and economic domains, along with rigorous evaluation protocols (Makridakis et al., 2020; 2022). Recent efforts focus on providing unified benchmarking frameworks and toolkits. TFB (Qiu et al., 2024), for instance, introduces a scalable suite covering multiple domains and methods, emphasizing reproducibility and systematic, aspect-based analysis. Furthermore, BasicTS (Liang et al., 2022) focuses on benchmarking multivariate time series forecasting, ProbTS (Zhang et al., 2024) benchmarks both point and distributional forecasting, and GIFT-Eval (Aksu et al., 2024) is designed for benchmarking foundation time-series models. Recently, ReC4TS (Liu et al., 2025b) benchmarks how different reasoning strategies enhance zero-shot time-series forecasting. In addition, Time-MMD (Liu et al., 2024a) and CiK (Ashok et al., 2024) are designed for benchmarking multimodal time-series forecasting. Different from all existing time-series benchmarks, our work uniquely focuses on evaluating the effectiveness of individual module-level design choices for time-series forecasting. To the best of our knowledge, this is the first benchmark to systematically address this underexplored yet critical aspect of model design.
@@
Auto-ML in Time-Series. Automated Machine Learning (AutoML) aims to automate the end-to-end process of applying machine learning, enhancing accessibility and efficiency. AutoML research in time series can be categorized as: (1) automated feature engineering, which generates relevant temporal features like lags and rolling statistics (Cerqueira et al., 2021; Costa, 2021); (2) automated model selection, searching across diverse model families from statistical methods (e.g., ARIMA) to machine learning (e.g., boosted trees) and deep learning architectures (e.g., LSTMs, Transformers) (Ying et al., 2020; Abdallah et al., 2022; Shchur et al., 2023); (3) hyperparameter optimization, using techniques like Bayesian optimization or evolutionary algorithms to tune model configurations (Wu et al., 2022; Fristiana et al., 2024); (4) neural architecture search, which specifically automates the design of deep learning model structures suitable for capturing complex temporal patterns (Rakhshani et al., 2020; Wu et al., 2023).